OpenStudy (anonymous):

what is the nature of solution of this equation x^2-2x+1=0

OpenStudy (anonymous):

2 non real solutions?

jhonyy9 (jhonyy9):

2 real

OpenStudy (anonymous):

calculate b^2-4ac 4-4*1*1 = 0 1 real solution

jhonyy9 (jhonyy9):

the same equales because the discriminant is zero

jhonyy9 (jhonyy9):

this equation can be wrote like (x-1) squared

jhonyy9 (jhonyy9):

so x=1 is solution

OpenStudy (anonymous):

\[(x-1)^{2}=0\] since it is a perfect square so it has real and equal roots

OpenStudy (anonymous):

(x-1)^2 = 0 x-1 = 0 x = 1

OpenStudy (anonymous):

one real solution